A typical incandescent bulb has an efficiency of about 10%. What current is needed for a bulb with a resistance of 1.5×10^2 Ω to output power of 6.0 W?


1
Expert's answer
2022-03-15T16:50:04-0400

Explanations & Calculations


  • Working power of the bulb

"\qquad\qquad
\begin{aligned}
\small P&=\small 6.0\,W\times\frac{100}{10}=60.0\,W
\end{aligned}"

  • The the current consumed at that power would be

"\qquad\qquad
\begin{aligned}
\small P&=\small I^2R\\
\small I &=\small \sqrt{\frac{60.0\,W}{1.5\times10^2\, \Omega}}= 6.3\times10^{-1}\,A
\end{aligned}"